# Weyl symmetry without the traceless condition

### v3

@@ II Local scale symmetry and the trace of the stress-energy tensor of matter
${\cal E}_{\mu\nu}\equiv G_{\mu\nu}-\frac{6}{\phi^{2}}\,\Theta^{(\phi)}_{\mu\nu},\;\Theta^{(\phi)}_{\mu\nu}\equiv\frac{1}{6}\left(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla^{2}\right)\phi^{2}-T^{(\phi)}_{\mu\nu}$. (14)
The modified SET of the scalar field $\Theta^{(\phi)}_{\mu\nu}$ ccj_ann_phys_1970 , has finite elements to all orders in the perturbation parameter and defines the same four-momentum and Lorentz generators as the conventional tensor: [^6] Note that in the expression for $\Theta^{(\phi)}_{\mu\nu}$ the standard stress-energy tensor of the scalar field $T^{(\phi)}_{\mu\nu}$, has the wrong sign. This is because, in the gravitational Lagrangian density (12), the kinetic energy density of the scalar field appears with the wrong sign, which is a requirement for the CGR gravitational Lagrangian density (12) to be invariant under WTs (11). This has no physical consequences, since $\phi$ is a gauge field that does not obey a specific EOM.
@@
Here we use the shorthand notation $\nabla^{2}\equiv g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$. If we take into account the trace of the tensor ${\cal E}_{\mu\nu}$:
@@
respectively. If we compare these equations, we obtain the following Ward identity:
@@
This identity is a consequence of the invariance of the CGR gravitational action (9) under CT (11). Actually, if this action is form-invariant under infinitesimal CTs:
@@
where for convenience (and only temporarily) we rescaled the conformal factor: $\Omega=e^{\theta}$, then $S_{\text{grav}}\rightarrow S_{\text{grav}}+\delta S_{\text{grav}}$, where $\delta S_{\text{grav}}=0$; so that, under (18), $S_{\text{grav}}\rightarrow S_{\text{grav}}$. Hence;
@@
where in the last equality, we substituted $\delta g^{\mu\nu}$ and $\delta\phi$ from infinitesimal CTs (18). Specifically, from the last equation, the Ward identity (17) follows.
After applying the variational principle of least action, from (13) we obtain the vacuum Einstein EOM and the Klein-Gordon-type EOM:
$\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}$ $=0\;\Rightarrow\;{\cal E}_{\mu\nu}=0\;\Rightarrow\;G_{\mu\nu}=\frac{6}{\phi^{2}}\,\Theta^{(\phi)}_{\mu\nu}$, (19)
$\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}$ $=0\;\Rightarrow\;\left(\nabla^{2}-\frac{1}{6}\,R\right)\phi=0,$ (20)
respectively. The trace of the Einstein EOM (19) reads $R+6\nabla^{2}\phi/\phi=0$, which, after some simple algebra, can be written as the KG-EOM (20). Hence, the latter is not an independent equation. Consequently, the scalar field $\phi$ does not obey an independent EOM. This is consistent with the new degree of freedom (DOF) associated with the Weyl symmetry of (9) since, in addition to the four DOFs for making spacetime diffeomorphisms, there is a DOF for making CTs (11), which is exhausted by the gauge field $\phi$.
Let us next consider matter fields coupled to CGR;
@@
where, since it is assumed in deser-1970 that the CTs do not transform the masses, the matter Lagrangian of time-like matter fields breaks LSS. In this case, the Ward identity for matter fields is not satisfied, so $\delta{\cal L}_{m}/\delta\phi=0$, because the matter Lagrangian density does not depend explicitly on $\phi$.
@@
respectively, where we have omitted divergence terms and assumed that,
@@
If we apply the variational principle, from (22) we get the following EOMs:
@@
Note that the Einstein EOM (19) has been modified by adding the SET of matter, whereas the KG-type EOM (20) is not modified since the Lagrangian density of the matter fields ${\cal L}_{m}=\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$ does not depend explicitly on $\phi$, that is, equation (23) takes place. This results in the following requirement. If we compare the trace of (24):
@@
with the EOM (25), we obtain the “traceless condition:” $T^{(m)}=0$. Therefore, only matter fields or fluids with vanishing SET trace: photons and radiation, in general, can be coupled to gravity in this local scale-invariant theory. [^7] Non-relativistic matter fields that behave like radiation can also be coupled to CGR faria-a .
@@
III Weyl symmetry does not require vanishing trace of the stress-energy tensor of matter
@@
In this section, we will demonstrate that, if we assume that the masses of timelike fields are themselves point-dependent fields that transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, while the energy density of a perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, it is not necessary to make any changes in (21) to get a consistent coupling of any matter fields. That is, under these assumptions, the coupling of the matter fields in (21) preserves the Weyl symmetry regardless of whether $T^{(m)}$ vanishes or not. Our demonstration is based on the Weyl invariance of matter action in this case, as demonstrated in quiros_prd_2025 for fundamental fields and perfect fluids.
@@
Under the above assumptions, for fermions and gauge bosons as well as for perfect fluids, the matter action (5) is not only invariant but also form-invariant under WTs. That is, under (7), the following equality takes place:
$\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})=\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{\partial}\hat{\chi}_{i},\hat{g}_{\mu\nu}).$ (27)
@@
From this equation it follows that if $g_{\mu\nu}$ is the physical metric with respect to the matter fields $\chi_{i}$; that is, the matter fields $\chi_{i}$ are minimally coupled to $g_{\mu\nu}$, the conformal metric $\hat{g}_{\mu\nu}$ is the physical metric with respect to the conformal matter fields $\hat{\chi}_{i}$ which are minimally coupled to $\hat{g}_{\mu\nu}$. In other words, if the metric $g_{\mu\nu}$ is the physical metric on some gauge, its conformal metric $\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$ is the physical metric on the conformal gauge. Equation (27) plays an important role because it shows that the total action (21) is Weyl invariant regardless of the type of matter coupled to gravity: fermions, gauge bosons, or perfect fluids.
@@
The transformation of masses $m\rightarrow\Omega^{-1}m$ and the consequent transformation of the energy density (and pressure) of a perfect fluid: $\rho\rightarrow\Omega^{-4}\rho$, is fundamental for the demonstration of form-invariance of the Lagrangian density of timelike fundamental matter fields quiros_prd_2025 . In this case, the Lagrangian density of matter fields is form-invariant under WTs, so the variational derivative: $\delta{\cal L}_{m}/\delta\phi$ obeys the Ward identity; that is, it is nonvanishing. This is the primary assumption underlying our approach to the Weyl symmetry of the CGR theory, which distinguishes our result from the well-known result that the matter SET trace must vanish if local scale invariance is a symmetry of the CGR theory.
Let us apply the variational procedure under the assumption that the action of the matter degrees of freedom coupled to gravity: fermions, gauge bosons, and perfect fluids, is form-invariant under the following WTs:
$g_{\mu\nu}\rightarrow\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\Omega^{-1}\phi,\;\chi_{i}\rightarrow\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\Omega^{-1}m.$ (28)
@@
Let us consider the Lagrangian density of matter fields: ${\cal L}_{m}={\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$. Since it is invariant under WTs (28), then, under infinitesimal Weyl transformations (18) plus $\delta\chi_{i}=w_{\chi_{i}}\theta\chi_{i}$ and $\delta m=-\theta m$, we have $\delta{\cal L}_{m}=\nabla_{\mu}V^{\mu}$, where $V^{\mu}$ is some vector or $\delta{\cal L}_{m}=0$. In any case, $\delta S_{m}=\int d^{4}x\,\delta{\cal L}_{m}=0$, so that under (18), $S_{m}\rightarrow S_{m}+\delta S_{m}=S_{m}$; that is, the matter action is invariant under infinitesimal Weyl transformations
@@
are satisfied. [^8] For classical gravitational problems, it is assumed that the matter fields are a priori given. From the last equality in (30), it follows the Ward identity:
@@
We have
@@
where in the last line of (34), in the factor multiplying $\delta\phi$ (expression within square brackets), we have taken into account the modified Ward identity (33). The variational principle leads to the following EOM:
@@
This is an outstanding result, since the trace of the matter SET is not required to vanish. Note that the KG-EOM (36) is not an independent equation, as it coincides with the trace of Einstein-type EOM (35). That is, the scalar field $\phi$ is not a dynamical field since it does not obey an independent EOM.
IV CGR: Physics beyond GR
is form-invariant under the WTs (28). The independent EOM derived from (37), is the following Einstein-type equation
${\cal E}_{\mu\nu}=\frac{6}{\phi^{2}}\,T^{(m)}_{\mu\nu}\;\Rightarrow\;G_{\mu\nu}$ $=\frac{6}{\phi^{2}}\left[\Theta^{(\phi)}_{\mu\nu}+T^{(m)}_{\mu\nu}\right$], (38)
@@ IV CGR: Physics beyond GR
Another independent equation is provided by the continuity equation. If we take the divergence of (38), and use the second Bianchi identity $\nabla^{\lambda}G_{\lambda\mu}=0$, and the identity $(\nabla_{\lambda}\nabla_{\mu}-\nabla_{\mu}\nabla_{\lambda})\nabla^{\lambda}\phi^{2}=R_{\mu\lambda}\nabla^{\lambda}\phi^{2}$, we obtain
@@
Then, if we substitute the trace of the Einstein EOM (36) in the latter equation, we get the following inhomogeneous continuity equation:
@@
The inhomogeneous term on the right-hand side (RHS) of (39) is required for the above equation to be form-invariant under WTs (11). This term represents a fifth force, which can be related to the point-dependent property of the mass parameter in the present theory $m(x)=\kappa\,\phi(x)$ bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 . Since timelike fields acquire masses through Yukawa interactions with the Higgs field $H$ after $SU(2)\times U(1)$ symmetry breaking, it is required to modify the standard model of particles (SMP) to get point-dependent masses. That is, the SMP must be Weyl invariant, and the $SU(2)\times U(1)$ symmetry breaking must not spoil Weyl symmetry. Therefore, the Higgs Lagrangian density must be appropriately modified bars-2014 ; mohamedi-2024 . In particular, the mass parameter $v$ in the Higgs potential, $V\propto(H^{\dagger}H-v^{2})^{2}$, must be replaced by a field: $v\rightarrow v_{0}\phi$, where $v_{0}$ is some dimensionless parameter bars-2014 . After the mentioned modification, the timelike fields acquire point-dependent masses through the Yukawa interaction. For example, for a fermion field, the Yukawa coupling is given by $\kappa_{f}\bar{\psi}\phi\psi=\bar{\psi}\,m_{f}\,\psi$ $\Rightarrow$ $m_{f}=\kappa_{f}\,\phi$.
Since $\phi$ is the gravitational coupling in CGR theory, that is, $\phi\sim M_{\text{pl}}$, a mass scale issue could arise. However, we will never be able to measure the point-dependent mass, $m(x)=\kappa\,\phi(x)$, in local experiments, because measuring sticks and clocks are made of atoms (matter) that suffer the same universal point-dependence, so the mass unit also varies: $u_{m}(x)=u_{0}\,\phi(x)$. In experiments, we measure the ratio faraoni_prd_2007
$\frac{m}{u_{m}}=\frac{\kappa}{u_{0}},$
that does not depend on the spacetime point. Hence, even if we assume $\phi_{0}\sim M_{\text{pl}}$, the value $\phi_{0}$ is not what is measured in experiments. In other words, the gigantic mass $m=\kappa\,\phi_{0}$ is compensated for by an equally gigantic mass unit $u_{m}=u_{0}\,\phi_{0}$, so there is no mass scale problem associated with point-dependent masses. The point-dependence of masses (and the related fifth force) can be observed only in “nonlocal” experiments; that is, in experiments that imply emission and absorption of photons coming from distant spacetime points; for example, a photon emitted by some atom located at some point $(t,\vec{x})$ is absorbed by an identical atom at the origin $(0,\vec{0})$. Since the mass of electrons is point-dependent, so is the energy of the emitted/absorbed photon. On the other hand, photon paths are not sensitive to WTs (the Riemannian null-geodesics are already Weyl invariant). This mismatch makes it possible to reveal the point-dependence of the masses, but adequate experiments must be designed (see Section VIII for further discussion of this subject).
IV.1 The fifth force
An interesting property of the CGR theory is related to the inhomogeneous continuity equation (39): $\nabla^{\lambda}T^{(m)}_{\lambda\mu}=T^{(m)}\partial_{\mu}\phi/\phi$. While this equation is satisfied by matter fluxes of stress and energy, the equivalent EOM for point-like particles is the following inhomogeneous EOM for time-like fields:
@@ IV.1 The fifth force
IV.2 Passive conformal transformations
It has been demonstrated in woodard-1986 (see also hobson-epjc-2022 ) that there is no physics beyond GR in CGR theory. The argument goes like this. To assess the physical consequences of CGR (in woodard-1986 called “conformal scalar-metric theory” in contrast to GR, called “purely metric theory”), we must rely on Weyl-invariant objects and operators. In GR theory, every physical question can be reduced to the VEV of some coordinate-invariant functional of the metric ${\cal F}[g]$. Similarly, every physical question in CGR can be reduced to the VEV of a coordinate and conformally invariant functional of $\phi$ and $g_{\mu\nu}$: $\hat{{\cal F}}[\phi,g]$. In woodard-1986 , it is demonstrated that the VEV of ${\cal F}[g]$ for the theory whose action is $S[g]$, equals the VEV of $\hat{{\cal F}}[\phi,g]$ for the theory whose action is $\hat{S}[\phi,g]$. Hence, any conformal scalar-metric theory is gauge equivalent to the corresponding pure metric theory.
Let us demonstrate the above result in an alternative way and show that the related point of view represents only one of two possible approaches to CTs (28): i) passive Weyl transformations (PWTs), and ii) active Weyl transformations (AWT’s) quiros_prd_2025 . Consider some abstract space of fields ${\cal M}_{f}$ where the metric $g_{\mu\nu}$, the gauge field $\phi$, the masses $m$, and any matter fields $\chi_{i}$ coupled to gravity, are “generalized coordinates.” Each point $P:(\phi,g_{\mu\nu},\chi_{i},m)$ in ${\cal M}_{f}$ represents a gravitational state. Passive CTs

### v4

@@ II Local scale symmetry and the trace of the stress-energy tensor of matter
${\cal E}_{\mu\nu}\equiv G_{\mu\nu}-\frac{6}{\phi^{2}}\,\Theta^{(\phi)}_{\mu\nu},\;\Theta^{(\phi)}_{\mu\nu}\equiv\frac{1}{6}\left(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla^{2}\right)\phi^{2}-T^{(\phi)}_{\mu\nu}$, (13)
where we use the shorthand notation $\nabla^{2}\equiv g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$. The modified SET of the scalar field $\Theta^{(\phi)}_{\mu\nu}$ ccj_ann_phys_1970 , has finite elements to all orders in the perturbation parameter and defines the same four-momentum and Lorentz generators as the conventional tensor: [^6] Note that in the expression for $\Theta^{(\phi)}_{\mu\nu}$ the standard stress-energy tensor of the scalar field $T^{(\phi)}_{\mu\nu}$, has the wrong sign. This is because, in the gravitational Lagrangian density (11), the kinetic energy density of the scalar field appears with the wrong sign, which is a requirement for the CGR gravitational Lagrangian density (11) to be invariant under CTs (10). This has no physical consequences, since $\phi$ is a gauge field that does not obey a specific EOM.
@@
If we take into account the trace of the tensor ${\cal E}_{\mu\nu}$:
@@
respectively. Therefore, the following Ward identity is obtained:
@@
This identity is a consequence of the invariance of the CGR gravitational action (8) under CTs (10). Actually, if this action is form-invariant under infinitesimal CTs:
@@
where for convenience (and only temporarily) we rescaled the conformal factor: $\Omega=e^{\theta}$, then $S_{\text{grav}}\rightarrow S_{\text{grav}}+\delta S_{\text{grav}}$, where $\delta S_{\text{grav}}=0$; so that, under (17), $S_{\text{grav}}\rightarrow S_{\text{grav}}$. We have;
@@
where in the last equality, we substituted $\delta g^{\mu\nu}$ and $\delta\phi$ from infinitesimal CTs (17). The Ward identity (16) follows from the last equation.
Let us consider matter fields coupled to CGR;
@@
where, in deser-1970 , local scale transformations (2) are considered rather than Weyl transformations (3). Hence, it is assumed in the mentioned bibliographic reference that the CTs do not transform the masses; that is, the matter Lagrangian of time-like matter fields breaks LSS. In this case, the Ward identity for matter fields is not satisfied, so $\delta{\cal L}_{m}/\delta\phi=0$, because the matter Lagrangian density does not depend explicitly on $\phi$.
@@
respectively, where we have omitted divergence terms and took into account that,
@@
From (19) we get the following EOMs:
@@
If we compare the trace of the Einstein EOM (21):
@@
with the Klein-Gordon (KG) type EOM (22), we obtain the “traceless condition:” $T^{(m)}=0$. Therefore, only matter fields and fluids with vanishing SET trace: photons and radiation, in general, can be coupled to gravity in this local scale-invariant theory. [^7] Non-relativistic matter fields that behave like radiation can also be coupled to CGR faria-a .
@@
III Conformal General Relativity with arbitrary matter fields
is form-invariant under the following Weyl transformations:
$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\hat{\phi}=\Omega^{-1}\phi,\;\chi_{i}\rightarrow\hat{\chi}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\hat{m}=\Omega^{-1}m.$ (25)
@@ III Conformal General Relativity with arbitrary matter fields
In this section, we will demonstrate that, if we assume that the masses of timelike fields are themselves point-dependent fields that transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, while the energy density of a perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, it is not necessary to make any changes in (18)/(24) to get a consistent coupling of any matter fields $\xi_{i}$ to Weyl invariant gravity. That is, under these assumptions, the coupling of the matter fields in (24) preserves Weyl symmetry, regardless of whether $T^{(m)}$ vanishes. Our demonstration is based on the Weyl invariance of matter action, as demonstrated in quiros_prd_2025 for fundamental fields and perfect fluids.
@@
Under the above assumptions, for fermions, gauge bosons, and perfect fluids, the matter action (4) is not only invariant but also form-invariant under WTs. That is, under (25), the following equality takes place:
${\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})={\cal L}_{m}(\hat{\chi}_{i},\hat{\partial}\hat{\chi}_{i},\hat{g}_{\mu\nu})\;\Leftrightarrow\;\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})=\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{\partial}\hat{\chi}_{i},\hat{g}_{\mu\nu}).$ (26)
@@
From this equation it follows that if $g_{\mu\nu}$ is the physical metric with respect to the matter fields $\chi_{i}$; that is, the matter fields $\chi_{i}$ are minimally coupled to $g_{\mu\nu}$, the conformal metric $\hat{g}_{\mu\nu}$ is the physical metric with respect to the conformal matter fields $\hat{\chi}_{i}$ which are minimally coupled to $\hat{g}_{\mu\nu}$. In other words, if the metric $g_{\mu\nu}$ is the physical metric on some gauge, its conformal metric $\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$ is the physical metric on the conformal gauge. Equation (26) plays an important role because it shows that the total Lagrangian density (24) is Weyl-invariant regardless of the type of matter coupled to gravity: fermions, gauge bosons, or perfect fluids.
@@
Given that the Lagrangian density of matter fields ${\cal L}_{m}$ is form-invariant under WTs (25), the variational derivative: $\delta{\cal L}_{m}/\delta\phi$ obeys the Ward identity; that is, it is nonvanishing. This is the primary assumption underlying our approach to the Weyl symmetry of the CGR theory, which distinguishes our result from the well-known result deser-1970 that the matter SET trace must vanish if local scale invariance is a symmetry of the CGR theory.
@@
Let us apply the variational procedure under the assumption that the action of the matter degrees of freedom coupled to gravity: fermions, gauge bosons, and perfect fluids, is form-invariant under WTs (25). We consider the Lagrangian density of matter fields: ${\cal L}_{m}={\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$. Since it is invariant under WTs (25), then, under infinitesimal Weyl transformations (17) plus $\delta\chi_{i}=w_{\chi_{i}}\theta\chi_{i}$ and $\delta m=-\theta m$, we have $\delta{\cal L}_{m}=0$, or $\delta{\cal L}_{m}=\nabla_{\mu}V^{\mu}$, where $V^{\mu}$ is some vector. In any case, $\delta S_{m}=\int d^{4}x\,\delta{\cal L}_{m}=0$, so that under (17), $S_{m}\rightarrow S_{m}+\delta S_{m}=S_{m}$; that is, the matter action is invariant under infinitesimal Weyl transformations
@@
are satisfied. [^8] For most of classical gravitational problems, it is assumed that the matter fields are a priori given. From the last equality in (28), it follows the Ward identity:
@@
Let us now apply the variational principle to derive the equations of motion. We have
@@
where in the last line of (32), in the factor multiplying $\delta\phi$ (expression within square brackets), we have taken into account the modified Ward identity (31). The variational principle leads to the following EOMs:
@@
This is an outstanding result because, while the EOM (33), (34), are form-invariant under Weyl transformations, the trace of the matter SET is not required to vanish. Note that the KG-EOM (34) is not an independent equation, as it coincides with the trace of Einstein-type EOM (33). That is, the scalar field $\phi$ is not a dynamical field, as it does not satisfy an independent EOM.
III.1 The “dark force”
The independent EOM derived from (24), is the Einstein-type equation (33);
${\cal E}_{\mu\nu}=\frac{6}{\phi^{2}}\,T^{(m)}_{\mu\nu}$, (35)
@@ III.1 The “dark force”
The continuity equation provides another independent equation. If we take the divergence of (35), and use the second Bianchi identity $\nabla^{\lambda}G_{\lambda\mu}=0$, and the identity $(\nabla_{\lambda}\nabla_{\mu}-\nabla_{\mu}\nabla_{\lambda})\nabla^{\lambda}\phi^{2}=R_{\mu\lambda}\nabla^{\lambda}\phi^{2}$, we obtain
@@
Then, if we substitute (34) in the latter equation, we get the following inhomogeneous continuity equation:
@@
The inhomogeneous term on the right-hand side (RHS) of (36) is required for the above equation to be form-invariant under WTs (25). This term represents a fifth force, which can be related to the point-dependent property of the mass parameter in the present theory: $m(x)=\kappa\,\phi(x)$ bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 . While (36) is satisfied by matter fluxes of stress and energy, the equivalent EOM for point-like particles is the following EOM for time-like fields:
@@
The assumption that $m=m(\phi)$ can have consequences for the standard model of particles (SMP). Actually, since timelike fields acquire masses through Yukawa interactions with the Higgs field $H$ after $SU(2)\times U(1)$ symmetry breaking, it is required to modify the SMP to get point-dependent masses. That is, the SMP must be Weyl-invariant, and the $SU(2)\times U(1)$ symmetry breaking must not spoil Weyl symmetry. Therefore, the Higgs Lagrangian density must be appropriately modified bars-2014 ; mohamedi-2024 . In particular, the mass parameter $v$ in the Higgs potential, $V\propto(H^{\dagger}H-v^{2})^{2}$, must be replaced by a field: $v\rightarrow v_{0}\phi$, where $v_{0}$ is some dimensionless parameter bars-2014 . After the mentioned modification, the timelike fields acquire point-dependent masses through the Yukawa interaction. For example, for a fermion field, the Yukawa coupling is given by $\kappa_{f}\bar{\psi}\phi\psi=\bar{\psi}\,m_{f}\,\psi$ $\Rightarrow$ $m_{f}=\kappa_{f}\,\phi$.
Notice that, as long as $\phi$ is the gravitational coupling in CGR theory, that is, $\phi\sim M_{\text{pl}}$, a mass scale issue could arise. However, we will never be able to measure the $\phi$-dependent mass, $m(x)=\kappa\,\phi(x)$, in local experiments, because measuring sticks and clocks are made of atoms (matter) that share the same universal $\phi$-dependence, so the mass unit also varies: $u_{m}(x)=u_{0}\,\phi(x)$. In experiments, we measure the ratio faraoni_prd_2007
$\frac{m}{u_{m}}=\frac{\kappa}{u_{0}},$
that does not depend on the spacetime point. Hence, even if we assume $\phi_{0}\sim M_{\text{pl}}$, the value $\phi_{0}$ is not what is measured in experiments. In other words, the gigantic mass $m=\kappa\,\phi_{0}$ is compensated for by an equally gigantic mass unit $u_{m}=u_{0}\,\phi_{0}$, so there is no mass scale problem associated with point-dependent masses. The point-dependence of masses (and the related fifth force) can be observed only in “nonlocal” experiments. For example, in experiments that imply emission and absorption of photons coming from distant spacetime points, where a photon emitted by some atom located at some point $(t,\vec{x})$ is absorbed by an identical atom at the origin $(0,\vec{0})$. Since the mass of electrons is point-dependent, so is the energy of the emitted/absorbed photon. On the other hand, photon paths are not sensitive to WTs (the Riemannian null-geodesics are already Weyl-invariant). This mismatch makes it possible to reveal the point-dependence of the masses, but adequate experiments must be designed (see Section VIII for further discussion of this subject).
IV CGR: Physics beyond GR
It has been demonstrated in hobson-epjc-2022 ; woodard-1986 that there is no physics beyond GR in CGR theory. [^10] In woodard-1986 CGR is named “conformal scalar-metric theory” in contrast to GR, named “purely metric theory.” While in woodard-1986 , it is shown that any “conformal scalar-metric theory” is gauge equivalent to the corresponding “pure metric theory”, in hobson-epjc-2022 , it is argued that if one works in terms of quantities that can be physically measured (in particular, the metric to which matter fields are minimally coupled), conformal Schwarzschild-de Sitter (SdS) metric does not lead to any different result than SdS metric; that is, “in any conformal frame, the trajectories followed by ordinary matter particles are merely the timelike geodesics of the SdS metric.”
Here we shall show that the above-cited result, although correct, is not universal. It represents only one of two possible approaches to WTs (25): i) passive Weyl transformations (PWTs), and ii) active Weyl transformations (AWT’s) quiros_prd_2025 . Let us consider some abstract space of fields ${\cal M}_{f}$ where the metric $g_{\mu\nu}$, the gauge field $\phi$, the masses $m$, and any matter fields $\chi_{i}$ coupled to gravity, are “generalized coordinates.” Each point $P:(\phi,g_{\mu\nu},\chi_{i},m)$ in ${\cal M}_{f}$ represents a “global gravitational state” (GGS). [^11] Here, by a global gravitational state we understand full knowledge of the fields at every spacetime point: $\phi=\phi(t,\vec{x})$, $g_{\mu\nu}=g_{\mu\nu}(t,\vec{x})$, $\chi_{i}=\chi_{i}(t,\vec{x})$ and $m=m(t,\vec{x})$. Passive Weyl transformations